OpenStudy (anonymous):

List all rational roots for these polynomial equations: x^3-x^2+2x-2=0 x^3=2x^2-8x-16=0 Find the roots of each polynomial equation: 2x^4-5x^3-17x^2+41x-21=0 A polynomial equation with rational coefficients has the given roots: find two additional roots. sqrt(5) and -sqrt(13) 2 + 3i and 6i (As in imaginary numbers) Find a third-degree polynomial equation with rational coefficients that has the given numbers as roots: 2 and -4i

OpenStudy (agent0smith):

Too many questions for one post :P x^3-x^2+2x-2=0 Notice x=1 will work... 1-1 + 2 - 2 = 0. So you can divide off that zero with synthetic division. You can also factor it by grouping: \[\Large x^2(x-1) + 2(x-1) = 0\]